Determine the specific gravity of a fluid having viscosity of 0.0a poise and kinematic viscosity of 0.0da stokes


1
Expert's answer
2021-08-29T16:57:17-0400

The density of a liquid can be expressed in terms of kinematic viscosity and viscosity:


"\rho_L=\frac{\mu_L}{\nu_L}."

Specific gravity is the ratio of density of a liquid to the density of water:


"s=\frac{\rho_L}{\rho_W}."

So:


"s=\frac{\mu_L}{\rho_W\nu_L}=\frac{0.01·10}{1000·0.01·10}=0.001."